OpenStudy (anonymous):

if x+1/x =8, find the x^2 + x^-2

OpenStudy (anonymous):

Is it \[{x+1 \over x}=8?\]

OpenStudy (anonymous):

no, it is x+(1/x)

OpenStudy (anonymous):

(x + 1/x)^2 = x^2 + 2 + 1/x^2

OpenStudy (anonymous):

It is further = x^2 + 2 + x^-2

OpenStudy (anonymous):

finally (x + 1/x)^2 - 2 = x^2 + x^-2 a

OpenStudy (anonymous):

\[x+{1 \over x}=8 \implies x^2+2+x^{-2}=16 \implies x^2+x^{-2}=14\]

OpenStudy (anonymous):

You hv x + 1/x = 8 Squaring both sides (x + 1/x)^2 = 8^2 = 16 S0, 16 - 2 = x^2 + x^-2 so 14 is the answer

OpenStudy (anonymous):

Sorry not 16, it's 64. So the final answer is 62.

OpenStudy (anonymous):

the answer I have are: a) 60, b)58, c)62, d)0, e)non So the answer must be e)

OpenStudy (anonymous):

Correct Anwar!!!

OpenStudy (anonymous):

So answer is 62

OpenStudy (anonymous):

option c)

OpenStudy (anonymous):

\[x+{1 \over x}=8 \implies (x+{1 \over x})^2=8^2 \implies x^2+2+x^{-2}=64 \] \[\implies x^2+x^{-2}=62\]

OpenStudy (anonymous):

Here is the corrected version You hv x + 1/x = 8 Squaring both sides (x + 1/x)^2 = 8^2 = 64 S0, 64 - 2 = x^2 + x^-2 so 62 is the answer

OpenStudy (anonymous):

You got the idea emunrradtvamg?

OpenStudy (anonymous):

yes, thanks guys its clear now

OpenStudy (anonymous):

You're welcome!